OpenStudy (anonymous):

in 1980 there were about 170 million vehicles and about 227 million people in the United States. The number of vehicles grows at 4% a year while the pop grows at 1% per year.when was there on average, one vehicle per person?

OpenStudy (wolf1728):

1980 veh = 170*(1.04)^years popul = 227*(1.01)^years So when does 170*(1.04)^years = 227*(1.01)^years ? taking logs of both sides years * 2.2304489214 + 0.0170333393 = years * 2.3560258572 + 0.0043213738 years * 2.2474822607 = years * 2.360347231 ******************************************************** Figured I'd try this by trial and error. In 1990 vehicles = 251.6415 people = 250.749

OpenStudy (wolf1728):

Seems as if I'm having problems taking the logarithm of a function of the type m * (n ^ years)

OpenStudy (wolf1728):

veh = 170*(1.04)^years popul = 227*(1.01)^years So, we make these equal to each other 170*(1.04)^years = 227*(1.01)^years 227/170 = (1.04)^years / (1.01)^years Taking logs of both sides log (227/170) = years * log (1.04/1.01) 0.1255769358 = years * (0.0127119655) years = 0.1255769358 / 0.0127119655 = 9.8786403767 years Yes, it's just that simple.